# Disentangling Knowledge Representations for Large Language Model Editing

### v1

@@ Abstract
Knowledge Editing has emerged as a promising solution for efficiently updating embedded knowledge in large language models (LLMs). While existing approaches demonstrate effectiveness in integrating new knowledge and preserving the original capabilities of LLMs, they fail to maintain fine-grained irrelevant knowledge — facts that share the same subject as edited knowledge but differ in relation and object. This challenge arises because subject representations inherently encode multiple attributes, causing the target and fine-grained irrelevant knowledge to become entangled in the representation space, and thus vulnerable to unintended alterations during editing. To address this, we propose DiKE, a novel approach that Disentangles Knowledge representations for LLM Editing (DiKE). DiKE consists of two key components: a Knowledge Representation Disentanglement (KRD) module that decomposes the subject representation into target-knowledge-related and -unrelated components, and a Disentanglement-based Knowledge Edit (DKE) module that updates only the target-related component while explicitly preserving the unrelated one. We further derive a closed-form, rank-one parameter update based on matrix theory to enable efficient and minimally invasive edits. To rigorously evaluate fine-grained irrelevant knowledge preservation, we construct FINE-KED, a new benchmark comprising fine-grained irrelevant knowledge at different levels of relational similarity to the edited knowledge. Extensive experiments across multiple LLMs demonstrate that DiKE substantially improves fine-grained irrelevant knowledge preservation while maintaining competitive general editing performance.
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention for their extensive knowledge storage and advanced reasoning capabilities (Zhao et al., 2024). However, the inherent noise in their training data and the continuous evolution of world knowledge often lead to inaccuracies and outdated information (Cao et al., 2021). To address these limitations, knowledge editing (Wang et al., 2024) has emerged as a promising solution, enabling precise and efficient updates to the knowledge embedded within LLMs. Among various approaches, parameter-modifying methods—which directly update the internal parameters of LLMs—are particularly appealing due to their ability to produce consistent outputs without requiring additional inference-time context or external memory. Representative methods include fine-tuning-based techniques (e.g., FT-L (Zhu et al., 2020)), meta-learning strategies (e.g., KE (Wang et al., 2025), MEND (Mitchell et al., 2022)), and locate-then-edit methods (e.g., ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025)). This work focuses specifically on parameter-modifying approaches.
@@
A successful knowledge editing method should not only accurately inject the desired updates but also minimize unintended interference with existing irrelevant knowledge. Recent research (Geva et al., 2023) suggests that LLMs retrieve stored knowledge by recalling facts associated with specific subjects. These retrieval processes center around subject representations, which encapsulate extensive attribute information related to that subject. Consequently, irrelevant knowledge can be classified into two categories based on semantic proximity to the subject of edited knowledge: fine-grained and coarse-grained irrelevant knowledge. Fine-grained irrelevant knowledge comprises facts that share the same subject as the target knowledge but differ entirely in relation and object, rendering them logically independent of the knowledge being edited. For example, when updating the knowledge from "The President of USA is Biden" to "The President of USA is Trump", a fine-grained irrelevant fact such as "The capital of USA is Washington" should remain unchanged despite sharing the subject USA. However, as they are encoded in the same subject representation, fine-grained irrelevant knowledge often becomes deeply entangled with the target knowledge within the model’s parameter and representation spaces (Hernandez et al., 2024; Qin et al., 2024), rendering it particularly susceptible to unintended alterations during editing. In contrast, coarse-grained irrelevant knowledge refers to facts that involve completely different subjects, relations, and objects, such as "Microsoft was founded by Bill Gates", and is typically distant enough from the target edit that inadvertent alteration is unlikely.
To conduct a preliminary analysis of how editing affects different types of irrelevant knowledge, we randomly sampled $1,000$ edit instances. For each, we selected one fine-grained and one coarse-grained irrelevant fact and evaluated the performance of existing editing methods, ROME (Meng et al., 2022), AlphaEdit (Fang et al., 2025), on LlaMA3(8B) in preserving them. As illustrated in Figure 1(a), the results show a marked discrepancy between the two categories, with fine-grained irrelevant knowledge being more vulnerable to unintended alterations. This observation highlights the inherent challenge of preserving fine-grained irrelevant knowledge during the editing process.
@@
To address the aforementioned challenge, we propose a novel locate-then-edit approach that Disentangles Knowledge representations for LLM Editing (DiKE) to effectively preserve fine-grained knowledge unrelated to the target edit, as depicted in Figure 2. Specifically, we first introduce a Knowledge Representation Disentanglement (KRD) module (§3.1) that disentangles the subject representation into target-knowledge-related and target-knowledge-unrelated components within the LLM’s representation space. Subsequently, to inject target knowledge into LLMs without impacting fine-grained irrelevant knowledge, we develop a Disentanglement-based Knowledge Edit (DKE) (§3.2) module. This module performs editing operations on the target-knowledge-related representation while constraining the target-knowledge-unrelated representation to remain unchanged. Furthermore, leveraging matrix theory, we derive a rank-one update formula that satisfies the these constraints, enabling the efficient update of model parameters. To comprehensively evaluate the effectiveness of our method in preserving fine-grained irrelevant knowledge, we construct a new dataset, FINE-KED4.1), which categorizes test instances into three levels based on the relational semantic similarity between the edited knowledge and its fine-grained irrelevant counterparts. Extensive experiments using GPT2-XL (1.5B), GPT-J (6B) and LLaMA-3 (8B) on FINE-KED and CounterFact demonstrate that DiKE effectively preserves fine-grained unrelated knowledge while achieving comparable general edit performance with other state-of-the-art editing methods.
We summarize our contributions as follows:
@@ 2.1 Autoregressive Language Model
where $\mathbf{W}_{in}^{l}$ and $\mathbf{W}_{out}^{l}$ are the parameter matrices of the first and second layers of the FFN, respectively, and $f(\cdot)$ is a non-linear activation function. For brevity, we omit the layer index $l$ in subsequent sections and denote $\mathbf{W}$ $=$ $\mathbf{W}_{out}^{l}$.
@@ 2.2 Knowledge Editing via Rank-one Updates
Knowledge Editing aims to modify or inject single or multiple pieces of knowledge into LLMs requiring full retraining. Each fact is typically represented as a triple $(s,r,o)$, where $s$ denotes subject, $r$ relation, and $o$ object, respectively. An edit sample is denoted as $e=(s,r,o,o^{*})$, representing the update of the original knowledge $(s,r,o)$ to a new knowledge $(s,r,o^{*})$.
Rank-one Model Editing, exemplified by methods such as ROME and MEMIT (Meng et al., 2022, 2023), which follow a locate-then-edit paradigm. These methods assume that factual knowledge is encoded in the feed-forward networks (FFNs) of the model as key-value pairs, where the first FFN layer generates keys and the second layer uses these keys to produce values. To inject new knowledge, the FFN parameters $\mathbf{W}$ are updated to $\mathbf{\hat{W}}$, such that the model associates a new key-value pair $(\mathbf{k}_{*},\mathbf{v}_{*})$. To avoid corrupting unrelated knowledge, an auxiliary set of key–value constraints $\mathbf{K}_{0}=[\mathbf{k}_{1};\mathbf{k}_{2};,\dots,;\mathbf{k}_{p}]$ and $\mathbf{V}_{0}=[\mathbf{v}_{1};\mathbf{v}_{2};,\dots,;\mathbf{v}_{p}]$ is introduced, representing knowledge to be preserved. The updated weights $\mathbf{\hat{W}}$ are obtained by solving the following objective:
[…]
The closed-form solution is given by:
@@ 3 Methodology
In this section, we present our proposed method, DiKE, with its overall architecture illustrated in Figure 2. The DiKE framework comprises two core components: (1) Knowledge Representation Disentanglement (KRD), which explicitly disentangles the subject representation into target-knowledge-related and -unrelated components within the LLM’s representation space, and (2) Disentanglement-based Knowledge Edit (DKE), which injects target knowledge into specific parameters of LLMs using the disentangled representations while preserving the fine-grained irrelevant knowledge.
Figure 2: Overview of DiKE architecture. The KRD module is first trained to disentangle target-knowledge-related and -unrelated representations. The DKE module then edits the related representation to inject new knowledge while constraining the unrelated part to preserve fine-grained irrelevant knowledge. A closed-form solution (Eq.(19)) is derived for parameter updates.(Note: The KRD module is pre-trained on the training set and does not require retraining for subsequent editing tasks)
@@ 3.1 Knowledge Representation Disentanglement
where $\mathbf{H}_{s}\in\mathbb{R}^{B\times d}$ represents $B$ negative subject representations sampled from the same batch. Here, $[\mathbf{z}_{e}^{u};\mathbf{H}_{s}]$ and $[\mathbf{z}_{e}^{r};\mathbf{H}_{s}]$ serve as the negative sample sets for $\mathbf{z}_{e}^{r}$ and $\mathbf{z}_{e}^{u}$, respectively. The InfoNCE loss is formulated as:
@@
Notably, the Knowledge Disentangler and Knowledge Recomposer, once trained, can be directly applied during the editing process without requiring retraining for each individual edit sample. Detailed training procedures are provided in Appendix G.
@@ 3.2 Disentanglement-based Knowledge Edit
In this section, we discuss how to leverage disentangled representations to perform precise parameter updates for injecting target knowledge, while explicitly preserving fine-grained irrelevant knowledge.
@@
$\mathbf{v}_{*}$ $=\mathbf{h}_{s}^{*}-\mathbf{h}_{s}^{p},\quad\text{where}\quad\mathbf{h}_{s}^{p}=\left(\mathbf{h}_{s}^{0}+\sum\nolimits_{l=1}^{L}\mathbf{a}_{s}^{l}+\sum\nolimits_{l=1}^{L-1}\mathbf{v}_{s}^{l}\right).$ (15)
@@
where $\mathbf{k}_{i}\in\mathbf{K}_{0}$, $\mathbf{h}_{s_{i}}$ and $\mathbf{h}_{r_{i}}$ denote the corresponding representations of subject $s_{i}$ and $\mathbf{r}_{i}$ in the edited layer. The full derivation is provided in Appendix D.1.
@@
$\hat{\mathbf{W}}=\underset{\hat{\mathbf{W}}}{\operatorname{argmin}}\bigg{(}$ $\underbrace{\left\|\hat{\mathbf{W}}\mathbf{k}_{*}-\mathbf{v}_{*}\right\|_{F}^{2}}_{\textbf{Target knowledge editing}}+\underbrace{\left\|\hat{\mathbf{W}}\mathbf{K}_{0}-\mathbf{V}_{0}\right\|_{F}^{2}}_{\textbf{Coarse-grained irrelevant knowledge preserving}}$ (18)
@@
where $\mathbf{E}$ is the identity matrix and $\mathbf{\Delta}_{\text{MEMIT}}$ represents the parameter update derived from MEMIT (Equation (4)). The detailed derivation is provided in Appendix D.2.
@@ 4 Experiments
In this section, we first introduce our newly constructed dataset, FINE-KED, designed to evaluate the impact of editing on fine-grained irrelevant knowledge. To provide a more comprehensive assessment of our DiKE’s performance in gereral editing scenarios, we further evaluate DiKE on the widely used CounterFact benchmark across GPT2-XL (1.5B), GPT-J (6B), and LLaMA3 (8B). More detailed experimental results are provided in Appendix H.
@@ 4.1 FINE-KED Dataset
FINE-KED is designed to evaluate the impact of editing methods on fine-grained irrelevant knowledge. For each edit sample $(s,r,o)$, we construct fine-grained irrelevant knowledge $(s,r^{\prime},o^{\prime})$ and categorize the dataset into three levels: Easy, Middle, and Hard, based on the semantic relatedness between relation $r$ and $r^{\prime}$ (examples of these levels are provided in Table 1). We employ Efficacy to quantify the success rate of edits and Relational Locality to assess the preservation of fine-grained irrelevant knowledge. Detailed information is provided in Appendix E.1.
@@ 4.2 Experimental Setups
Datasets and Evaluation Metrics. In addition to FINE-KED, we evaluate our method on COUNTERFACT (Meng et al., 2022), which is designed to assess the insertion of counterfactual knowledge into models. Performance on this dataset is measured using three key metrics: Efficacy Score, Paraphrase Score, and Neighborhood Score. Further details are shown in Appendix E.2.
Baselines. Our experiments are conducted on three LLMs: GPT2-XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021) and LLaMA3 (8B) (Dubey et al., 2024). We compare our method with a number of knowledge editing methods: Fine-Tuning (FT) (Zhu et al., 2020), MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), MEMIT(Meng et al., 2023), and AlphaEdit (Fang et al., 2025). To further validate the superiority of DiKE in preserving the fine-grained irrelevant knowledge, we compare it with two variant models ROME-C and MEMIT-C on FINE-KED, which directly incorporate multiple additional relational constraints $(s,r_{i},o_{i})$ into the LLM in each editing. The implementation details of baselines and DiKE in Appendix G.

### v2

@@ Abstract
Knowledge Editing has emerged as a promising solution for efficiently updating embedded knowledge in large language models (LLMs). While existing approaches demonstrate effectiveness in integrating new knowledge and preserving the original capabilities of LLMs, they fail to maintain fine-grained irrelevant knowledge, namely facts that share the same subject as edited knowledge but differ in relation and object. This challenge arises because subject representations inherently encode multiple attributes, causing the target and fine-grained irrelevant knowledge to become entangled in the representation space, and thus vulnerable to unintended alterations during editing. To address this, we propose DiKE, a novel approach that Disentangles Knowledge representations for LLM Editing (DiKE). DiKE consists of two key components: a Knowledge Representation Disentanglement (KRD) module that decomposes the subject representation into target-knowledge-related and -unrelated components, and a Disentanglement-based Knowledge Edit (DKE) module that updates only the target-related component while explicitly preserving the unrelated one. We further derive a closed-form, rank-one parameter update based on matrix theory to enable efficient and minimally invasive edits. To rigorously evaluate fine-grained irrelevant knowledge preservation, we construct FINE-KED, a new benchmark comprising fine-grained irrelevant knowledge at different levels of relational similarity to the edited knowledge. Extensive experiments across multiple LLMs demonstrate that DiKE substantially improves fine-grained irrelevant knowledge preservation while maintaining competitive general editing performance.
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention for their extensive knowledge storage and advanced reasoning capabilities (Zhao et al., 2024). However, the inherent noise in their training data and the continuous evolution of world knowledge often lead to inaccuracies and outdated information (Cao et al., 2021). To address these limitations, knowledge editing (Wang et al., 2024) has emerged as a promising solution, enabling precise and efficient updates to the knowledge embedded within LLMs. Among various approaches, parameter-modifying methods that directly update the internal parameters of LLMs are particularly appealing due to their ability to produce consistent outputs without requiring additional inference-time context or external memory. Representative methods include fine-tuning-based techniques (e.g., FT-L (Zhu et al., 2020)), meta-learning strategies (e.g., KE (Wang et al., 2025), MEND (Mitchell et al., 2022)), and locate-then-edit methods (e.g., ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025)). This work focuses specifically on parameter-modifying approaches.
@@
A successful knowledge editing method should not only accurately inject the desired updates but also minimize unintended interference with existing irrelevant knowledge. Recent research (Geva et al., 2023) suggests that LLMs retrieve stored knowledge by recalling facts associated with specific subjects. These retrieval processes center around subject representations, which encapsulate extensive attribute information related to that subject. Consequently, irrelevant knowledge can be classified into two categories based on semantic proximity to the subject of edited knowledge: fine-grained and coarse-grained irrelevant knowledge. Fine-grained irrelevant knowledge comprises facts that share the same subject as the target knowledge but differ entirely in relation and object, rendering them logically independent of the knowledge being edited. For example, when updating the knowledge from ”The President of USA is Biden” to ”The President of USA is Trump”, a fine-grained irrelevant fact such as ”The capital of USA is Washington” should remain unchanged despite sharing the subject USA. However, as they are encoded in the same subject representation, fine-grained irrelevant knowledge often becomes deeply entangled with the target knowledge within the model’s parameter and representation spaces (Hernandez et al., 2024; Qin et al., 2024), rendering it particularly susceptible to unintended alterations during editing. In contrast, coarse-grained irrelevant knowledge refers to facts that involve completely different subjects, relations, and objects, such as ”Microsoft was founded by Bill Gates”, and is typically distant enough from the target edit that inadvertent alteration is unlikely.
To conduct a preliminary analysis of how editing affects different types of irrelevant knowledge, we randomly sampled $1,000$ edit instances. For each, we selected one fine-grained and one coarse-grained irrelevant fact and evaluated the performance of existing editing methods, FT-L (Zhu et al., 2020), MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), and AlphaEdit (Fang et al., 2025), on LlaMA3(8B) in preserving them. As illustrated in Figure 1(a), the results show a marked discrepancy between the two categories, with fine-grained irrelevant knowledge being more vulnerable to unintended alterations. This observation highlights the inherent challenge of preserving fine-grained irrelevant knowledge during the editing process.
@@
To address the aforementioned challenge, we propose a novel locate-then-edit approach that Disentangles Knowledge representations for LLM Editing (DiKE) to effectively preserve fine-grained knowledge unrelated to the target edit, as depicted in Figure 2. Specifically, we first introduce a Knowledge Representation Disentanglement (KRD) module (§3.1) that disentangles the subject representation into target-knowledge-related and target-knowledge-unrelated components within the LLM’s representation space. Subsequently, to inject target knowledge into LLMs without impacting fine-grained irrelevant knowledge, we develop a Disentanglement-based Knowledge Edit (DKE) (§3.2) module. This module performs editing operations on the target-knowledge-related representation while constraining the target-knowledge-unrelated representation to remain unchanged. Furthermore, leveraging matrix theory, we derive a rank-one update formula that satisfies the these constraints, enabling the efficient update of model parameters. To comprehensively evaluate the effectiveness of our method in preserving fine-grained irrelevant knowledge, we construct a new dataset, FINE-KED1), which categorizes test instances into three levels based on the relational semantic similarity between the edited knowledge and its fine-grained irrelevant counterparts. Extensive experiments using GPT2-XL (1.5B), GPT-J (6B) and LLaMA-3 (8B) on FINE-KED and CounterFact demonstrate that DiKE effectively preserves fine-grained unrelated knowledge while achieving comparable general edit performance with other state-of-the-art editing methods. We summarize our contributions as follows:
@@ 2.1 Autoregressive Language Model
where $\mathbf{W}_{in}^{l}$ and $\mathbf{W}_{out}^{l}$ are the parameter matrices of the first and second layers of the FFN, respectively, and $f(\cdot)$ is a non-linear activation function. For brevity, we omit the layer index $l$ in subsequent sections and denote $\mathbf{W}$ == $\mathbf{W}_{out}^{l}$.
@@ 2.2 Knowledge Editing via Rank-one Updates
Knowledge Editing aims to modify or inject single or multiple pieces of knowledge into LLMs without requiring full retraining. Each fact is typically represented as a triple $(s,r,o)$, where $s$ denotes subject, $r$ relation, and $o$ object, respectively. An edit sample is denoted as $e=(s,r,o,o^{*})$, representing the update of the original knowledge $(s,r,o)$ to a new knowledge $(s,r,o^{*})$.
Rank-one Knowledge Editing, exemplified by methods such as ROME and MEMIT (Meng et al., 2022; 2023), which follow a locate-then-edit paradigm. These methods assume that factual knowledge is encoded in the feed-forward networks (FFNs) of the model as key-value pairs, where the first FFN layer generates keys and the second layer uses these keys to produce values. To inject new knowledge, the FFN parameters $\mathbf{W}$ are updated to $\mathbf{\hat{W}}$, such that the model associates a new key-value pair $(\mathbf{k}_{*},\mathbf{v}_{*})$. To avoid corrupting unrelated knowledge, an auxiliary set of key–value constraints $\mathbf{K}_{0}=[\mathbf{k}_{1};\mathbf{k}_{2};,\dots,;\mathbf{k}_{p}]$ and $\mathbf{V}_{0}=[\mathbf{v}_{1};\mathbf{v}_{2};,\dots,;\mathbf{v}_{p}]$ is introduced, representing knowledge to be preserved.
Adopting the least-squares formulation from MEMIT (Meng et al., 2023), the updated weights $\mathbf{\hat{W}}$ are obtained by solving the following objective:
$\mathbf{\mathbf{\hat{W}}}=\underset{\mathbf{\hat{W}}}{\operatorname{argmin}}(\|\mathbf{\hat{W}}\mathbf{k}_{*}-\mathbf{v}_{*}\|_{F}^{2}+\|\mathbf{\hat{W}}\mathbf{K}_{0}-\mathbf{V}_{0}\|_{F}^{2}).$ (3)
The closed-form solution derived via the normal equation as in MEMIT, is given by:
@@ 3 Methodology
In this section, we present our proposed method, DiKE, with its overall architecture illustrated in Figure 2. The DiKE framework comprises two core components: (1) Knowledge Representation Disentanglement (KRD), which explicitly disentangles the subject representation into target-knowledge-related and -unrelated components within the LLM’s representation space, and (2) Disentanglement-based Knowledge Edit (DKE), which injects target knowledge into specific parameters using the disentangled representations while preserving the fine-grained irrelevant knowledge.
Figure 2: Overview of the DiKE architecture. The framework operates in two distinct phases. (Left) KRD Training: The module extracts subject and relation representations from the LLM and learns to disentangle them into target-knowledge-related and -unrelated components via optimizing disentanglement, constraint, and reconstruction objectives. (Right) DKE Editing: During the editing phase, the pre-trained Disentangler is frozen. The DKE module utilizes the disentangled representations to derive a closed-form rank-one parameter update (Eq. (19)), which injects new knowledge into the target-related component while explicitly constraining the unrelated component to preserve fine-grained irrelevant knowledge. Note that the KRD module is pre-trained on the training set and does not require retraining for subsequent editing tasks.
@@ 3.1 Knowledge Representation Disentanglement
where $\mathbf{H}_{s}\in\mathbb{R}^{B\times d}$ represents $B$ negative subject representations sampled from the same batch. Here, $\mathbf{z}_{e}^{u};\mathbf{H}_{s}]$ and $\mathbf{z}_{e}^{r};\mathbf{H}_{s}]$ serve as the negative sample sets for $\mathbf{z}_{e}^{r}$ and $\mathbf{z}_{e}^{u}$, respectively. The InfoNCE loss is formulated as:
@@
Notably, the Knowledge Disentangler and Knowledge Recomposer, once trained, can be directly applied during the editing process without requiring retraining for each individual edit sample. Detailed training procedures are provided in Appendix F.1.
@@ 3.2 Disentanglement-based Knowledge Edit
In this section, we discuss how to leverage disentangled representations to perform parameter updates for injecting target knowledge, while explicitly preserving fine-grained irrelevant knowledge.
@@
$\mathbf{v}_{*}$ $=\mathbf{h}_{s}^{*}-\mathbf{h}_{s}^{p},\quad\text{where}\quad\mathbf{h}_{s}^{p}=\mathbf{h}_{s}^{0}+\sum\nolimits_{l=1}^{L}\mathbf{a}_{s}^{l}+\sum\nolimits_{l=1}^{L-1}\mathbf{v}_{s}^{l}.$ (15)
@@
where $\mathbf{k}_{i}\in\mathbf{K}_{0}$, $\mathbf{h}_{s_{i}}$ and $\mathbf{h}_{r_{i}}$ denote the corresponding representations of subject $s_{i}$ and $\mathbf{r}_{i}$ in the edited layer. The full derivation is provided in Appendix C.1.
@@
[…]
@@
where $\mathbf{E}$ is the identity matrix and $\mathbf{\Delta}_{\text{MEMIT}}$ represents the parameter update derived from MEMIT (Equation (4)). The detailed derivation is provided in Appendix C.2.
@@ 4 Experiments
In this section, we first introduce our newly constructed dataset, FINE-KED, designed to evaluate the impact of editing on fine-grained irrelevant knowledge. To provide a more comprehensive assessment of our DiKE’s performance in gereral editing scenarios, we further evaluate DiKE on the widely used CounterFact and Mquake-3K benchmarks across GPT2-XL (1.5B), GPT-J (6B), and LLaMA3 (8B). More detailed experimental results are provided in Appendix G.
@@ 4.1 FINE-KED Dataset
FINE-KED is designed to evaluate the impact of editing methods on fine-grained irrelevant knowledge. For each edit sample $(s,r,o)$, we construct fine-grained irrelevant knowledge $(s,r^{\prime},o^{\prime})$ and categorize the dataset into three levels: Easy, Middle, and Hard, based on the semantic relatedness between relation $r$ and $r^{\prime}$ (examples of these levels are provided in Table 1). We employ Efficacy to quantify the success rate of edits and Relational Locality to assess the preservation of fine-grained irrelevant knowledge. Detailed information is provided in Appendix D.1.
@@ 4.2 Experimental Setups
Datasets and Evaluation Metrics. In addition to FINE-KED, we conduct experiments on COUNTERFACT (Meng et al., 2022), a benchmark designed to evaluate the insertion of counterfactual knowledge into LLMs. Performance is assessed using three key metrics: Efficacy Score, Paraphrase Score, and Neighborhood Score. Further details are in Appendix D.2. To further investigate DiKE’s generalization capability in more complex scenarios, we also evaluate on the multi-hop reasoning dataset MQuAKE-3K Zhong et al. (2023). Dataset descriptions are provided in Appendix D.3, with corresponding results reported in Appendix G.2.
Baselines. Our experiments are conducted on three LLMs: GPT2-XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021) and LLaMA3 (8B) (Dubey et al., 2024). We compare our method with a number of knowledge editing methods: Fine-Tuning (FT) (Zhu et al., 2020), MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), and AlphaEdit (Fang et al., 2025). To further validate the superiority of DiKE in preserving the fine-grained irrelevant knowledge, we compare it with two variant models ROME-C and MEMIT-C on FINE-KED, which directly incorporate multiple additional relational constraints $(s,r_{i},o_{i})$ into the LLM in each editing. The implementation details of baselines and DiKE in Appendix F.1.